# Super resonance: Breaking the bandwidth limit of resonant modes and its application to flow control

### v1

@@ Abstract
We report the discovery of super resonance—a new regime of resonant behavior in which a mode’s out-of-phase response persists far beyond its classical bandwidth. This effect emerges from a coiled phononic structure composed of a locally resonant elastic metamaterial and architected to support multiple internal energy pathways. These pathways converge at a single structural point, enabling extended modal dominance and significantly broadening the frequency range over which a resonant phase is sustained. We demonstrate by direct numerical simulations the implications of this mechanism in the context of flow instability control, where current approaches are inherently constrained by the characteristically narrow spectral bandwidth of conventional resonances. Using a super-resonant phononic subsurface structure interfacing with a channel flow, we show passive simultaneous suppression of four unstable flow perturbations across a frequency range more than five times wider than that is achievable with a standard resonance in an equivalent uncoiled structure. By enabling broadband, passive control of flow instabilities, super resonance overcomes a longstanding limitation in laminar flow control strategies. More broadly, it introduces a powerful new tool for phase-engineered wave–matter interaction. The ability to preserve out-of-phase modal response across wide spectral ranges establishes a fundamental advance in the physics of resonance, with far-reaching implications for suppressing fully developed turbulent flows and beyond.
@@ INTRODUCTION
The harmonic oscillator is one of the most fundamental constructs in physics, serving as a building block across classical mechanics, quantum theory, solid-state physics, and wave dynamics [1, 2, 3, 4, 5]. Its resonance behavior—marked by usually a sharply peaked amplitude response and a narrow out-of-phase spectral region that follows—underpins our understanding of modal systems, from phonons and photons to acoustic and electromagnetic cavities. In both isolated and coupled forms, harmonic oscillator modes are intrinsically limited in bandwidth; for coupled systems, the phase response is tightly overridden by the presence of nearby resonances. This long-standing narrowband constraint has limited the scope of resonance-based control strategies across physics and engineering. In this work, we introduce super resonance: a new regime of modal behavior in which a structural mode sustains an out-of-phase response that persists far beyond its classical resonance bandwidth, despite the nominal presence of adjacent resonant peaks. This behavior is achieved through a uniquely architected coiled phononic structure featuring multiple internal elastic energy pathways that converge to a single interface point. This configuration enables extended modal phase persistence, laying the foundation for broadband control in systems traditionally constrained by narrowband resonance dynamics—including the passive suppression of multiple unstable flow perturbations, which is the focus of this investigation.
@@
To overcome these limitations, the concept of phononic subsurfaces (PSubs) has been introduced as a passive intervention approach capable of precise and responsive control of flow instabilities [19]. A PSub comprises a finite phononic structure that is placed beneath the fluid-structure interface. Given its finite size, it represents a truncated [20, 21, 22] phononic material [23, 24, 25, 26], such as a phononic crystal that exhibits Bragg scattering [19, 27, 28, 29, 30], or a locally resonant elastic metamaterial that exhibits resonance hybridizations [28, 31]. The unit-cell and finite-structure characteristics of a PSub may be designed to passively force incident flow instability waves to encounter anti-resonance conditions. This causes a decrease, near the surface, in the production rate of the kinetic energy of the continuously incoming flow perturbation waves, and subsequently their attenuation across the region covered by the PSub. The outcome is a local reduction in skin-friction drag where the PSub unit is installed. Recent developments extended the concept to downstream control, by either adopting a multi-input, multi-output configuration [32], employing a lattice of PSubs [33], or using PSubs tuned to intermediate phase values [34]. The tuning of a single PSub structure depends on knowledge of the frequency, phase, and wavenumber of the relevant flow instabilities, including incidents involving multiple coexisting modes.
@@
While this passive phonon-enabled approach has shown great promise, in areas ranging from transition delay [32, 33] to control of unsteady flows [35], and recently extending to disturbance attenuation in hypersonic flows [34], its effectiveness remains limited by the characteristically narrowband nature of the modal response in the interfacing structure. This limitation also applies to other conventional resonators such as Helmholtz resonators [36]. Extensive investigations have been done on various PSub configurations, but due to this constraint are all confined to spectrally narrow phase response—whether, for example, due to a truncation resonance  [^1] A band-gap truncation resonance may exhibit a modestly broader out-of-phase bandwidth than a standard structural resonance—particularly when the band gap is wide [28]. Yet it remains fundamentally a conventional resonance subject to the same narrowband limitations described above. located within a band gap [19, 27, 29] or a subwavelength structural resonance influenced by the presence of local resonators [28, 31]. This narrowband restriction reduces the applicability of PSubs to realistic settings, such as aircraft operation, where flow instabilities can emerge over a relatively broad frequency range  [^2] These instability bands depend on the Reynolds number $Re$ and can be predicted a priori through linear stability analysis [70, 71, 27]. In this paper, by introducing super resonance, we present a new PSub concept that facilitates a significant broadening of the frequency range of operation.
@@
Numerous studies across physics and engineering have investigated different forms of enhanced resonant behavior, often highlighting superior properties. Examples include acoustic super-scattering from elastic shells [39], strong plasmonic field enhancement and lasing [40], high-Q supercavity modes enabled by bound states in the continuum in subwavelength dielectric resonators [41], and intensified Mie resonances in dielectric microspheres [42]. In each case, the response is intensified through distinct means—boundary-induced coherence, material gain, non-radiating interference effects, or structural symmetry, respectively. However, the underlying mechanism remains rooted in conventional resonant behavior: a narrowband spectral peak that fundamentally constrains modal reach. In contrast, the current work targets broadening the phase bandwidth of a target resonance beyond its natural limits. The core premise is to intervene in key energy pathways that govern the mechanistic formation of the resonance.
@@
In any resonant system, a peak response arises due to an excitation at a specific spatial location (or domain), and, similarly, there is a distinct spatial location (or domain) where the response may be probed. For a simple harmonic oscillator, there is only a single point of excitation and a single response point, creating a resonant peak with its characteristic narrow bandwidth. The same characteristic narrow band holds for modes in a coupled oscillator system. Here we consider a multi-degree-of-freedom system composed of coupled oscillators. But we reconfigure the spatial topology to allow multiple pathways for the energy to converge towards a single response point in space, selected to coincide with the excitation point. In this manner, the modal response of the system exhibits a reconstructed mode with a significantly widened bandwidth, exceeding the fundamental bandwidth associated with a conventional resonant mode.
@@
For application to the flow stabilization problem, the concept is realized in the form of a PSub architecture composed of a coiled beam structure, configured such that more than one location along the structure interfaces with a single point in the flow. The coiling is done by a series of full $180^{\circ}$ turns, incorporating rotational locking at the turning junctions-an approach that fully preserves the phonon band structure [43]. By following this design strategy, we observe remarkable broadening of a performance metric that governs the strength and spectral extent of intervention with the flow [19, 28, 31].
@@
Figure 1: Coiled PSub design: A schematic of the concept of a coiled phononic subsurface with multiple structural connectivity. The physical model, shown in (a), features a central voided beam that is rotationally locked at the turning junctions, leading to a full preservation of the phonon band structure [43]. The beam “trunk" has an array of branching resonators in the form of small cantilevered beams-this configuration in its entirety is amenable to fabrication. The dimensions of the PSub unit cell are as follows based on the parameters shown in (a): $L_{\text{UC}}=20$ $m$, $w_{\text{c}}=1.44$ $m$, $h_{\text{c}}=4.63$ $m$, $l_{\text{h}}=1.17$ $m$, $w_{\text{h}}=1.12$ $m$, $l_{1}=18.23$ $m$, $w_{1}=2.06$ $m$, $l_{2}=12.93$ $m$, $w_{2}=4.68$ $m$, $w_{\text{f}}=0.4$ $m$. The central beam component is composed of ABS polymer with density $\rho_{\text{ABS}}=1200$ kg/m3 and Young’s modulus $E_{\text{ABS}}=1$ GPa. The resonant branches are made from aluminium with density $\rho_{\text{Al}}=2700$ kg/m3 and Young’s modulus $E_{\text{Al}}=68.8$ GPa. In our computational investigation, this 3D structure is modeled in the form of a 1D rod, as shown in (b), with effective Young’s modulus $E_{\text{eff}}=0.2$ GPa and density $\rho_{\text{eff}}=450$ kg/m3, containing an array of spring-mass resonators attached based on the following parameters: $L_{\text{UC}}=20$ $m$, $k_{\text{res},1}=2.6\times 10^{10}$ N/m, $k_{\text{res},2}=7.11\times 10^{11}$ N/m, $m_{\text{res},1}=1.62$ g, $m_{\text{res},2}=4.2$ g. Material damping is introduced to the rod component in the form of viscous proportional damping with constants $q_{1}=0$ and $q_{2}=6\times 10^{-8}$ [28]. Given that our interest is in only the longitudinal motion, the rod model provides a good approximation of the coiled PSub response, owing to the rotational locking applied at the junctions. The selected simple model parameters deviate less than 2% from the actual unit-cell equivalent configuration. With one coiling cycle, we have two structural points that are simultaneously interfacing with the flow, labeled Junctions 1 and 3. With two coiling cycles, Junctions 1, 3, and 5 are in contact with the flow, and so on in an odd-number progression. We find that with three coiling cycles, comprising Junctions 1, 3, 5, 7, super-resonant behavior emerges (see Fig. 2). While the coiled PSub design is confined to a specific finite dimension along the spanwise direction, its response can still be considered one-dimensional when normalized by the surface area interacting with the flow, denoted by AA. In this work, we consider only a single coiled PSub covering the full span of the channel. However, application to longer spanwise distances (e.g., for wider channels or boundary-layer flows) may be realized by installation of a periodic arrangement of the coiled PSub along that direction (side view). Additional rows may also be added along the streamwise direction to form a PSub lattice setup [33].
@@
Figure 2: Evidence of super resonance: Coiled PSub dispersion and vibration characteristics: (a) Dispersion, (b-c) harmonic frequency response, and (d) performance metric for the coiled PSub with multiple structural connectivity. The dispersion is unchanged with the number of coiling cycles since the design of the unit cell is the same in all cases and band structure preservation is maintained due that rotational lockings [43]. The amplitude and phase frequency dependency for the uncoiled (0 coils) and 3-cycle coiled configurations are shown in (b-c). Positive (red) and negative (green) regions of $P$ represent destabilization and stabilization, respectively, upon employment of the PSub for flow control. Zero coiling cycles represent a PSub with 1 point-of-contact with the flow, while 3 coiling cycles represents the same PSub but with 4 points-of-contact with the flow, as demonstraed in Fig. 1. As the number of coils is increased from 1 to 3, the destabilization and stabilization regions across the frequency domain are narrowed and broadened, respectively. With 3 coiling cycles, super resonance emerges yielding a perfectly contiguous stabilization band that is nearly two times larger than that of the uncoiled reference case. A quasi-super resonance regime also emerges, one that is effectively contiguous and nearly four times larger than that of the reference configuration.
@@
In its nominal uncoiled form, the PSub considered comprises a finite linear elastic metamaterial consisting of 10 unit cells, where each is formed from three central voided beam-like segments with two distinct branching cantilever beams in-between, serving as local resonators; see illustration with geometric details in Fig. 1. To make extensive coupled fluid-structure simulations feasible, this phononic structure is modeled as an elastic rod with two mass-spring resonators attached per unit cell, similar to the configuration of Ref. [28]. The underlying reduce-order-modeling is facilitated by a combined homogenization and machine learning-algorithm [53, 54]; see Appendix for details. The PSub is free to deform at the edge interfacing with the flow (top) as well as all other intermediate locations throughout its structure, but is fixed at the other end (bottom). The PSub elastic domain is allowed to deform in complete independence from the adjacent rigid wall, thus the flow “experiences" a deformation in the form of a uniform profile across a given fluid-PSub interface junction [19, 28]. The length of the unit cell along the wall-normal direction is $L_{\mathrm{UC}}=2$ $\mathrm{cm}$ (i.e., total uncoiled PSub length is 2020 $\mathrm{cm}$). The resonator frequencies are set to $\Omega_{\mathrm{res,1}}=5000$ $\mathrm{Hz}$ and $\Omega_{\mathrm{res,2}}=20,000$ $\mathrm{Hz}$ by tuning their point masses to be three and five times heavier than the total mass of the unit-cell base, i.e., $m_{\mathrm{res,1}}=3\times\rho L_{\mathrm{UC}}$ and $m_{\mathrm{res,2}}=5\times\rho L_{\mathrm{UC}}$, where $\rho$ is the base material effective density. Hence, the stiffnesses of the resonator springs are $k_{\mathrm{res,n}}=m_{\mathrm{res,n}}(2\pi f_{\mathrm{res,n}})^{2}$, where $n=1,2$.
@@
 The PSub coiling protocol is done as follows. A full $180^{\circ}$ turn is applied to each unit cell, starting with the second one, and up to twice the number of coiling cycles. Thus, with $n$ coiling cycles, $2n$ full turns are applied at the intersections of the first $2n+1$ unit cells. In order to guarantee that the phonon band structure is preserved, rotationally rigid junctions are placed at these intersections to lock the rotation degree of freedom [43]. On the top side, these junctions converge to a $n+1$ points in contact with the flow, with the ability to interact independently with it. At these points, the flow represents the excitation source from the view point of the coiled PSub. Since these points are confined within a small control region along the streamwise direction compared to the perturbation wavelength(s), from the flow point of view they are effectively perceived as a single point. This means that the rotationally rigid junctions transmit and distribute the total force equally from the flow to the coiled PSub, and the coiled PSub’s response is the cumulative contribution of all the points in contact with the flow. See Fig. 1 for a schematic and geometric details.
@@
Figure 3: Coiled PSub performance by DNS: (a) Stability map in the frequency-Reynolds number domain obtained by solving the Orr-Sommerfeld equation [45, 46] for the channel. (b) Zoom-in of the performance metric plot for the super-resonant 3-cycle coiled PSub, with frequencies of the four unstable TS waves marked. (c) Time-averaged perturbation kinetic energy as a function of streamwise position in the channel. The results of five separate simulations are shown: four in which a distinct TS instability wave is controlled (left axis), and one where all four TS instability waves are controlled simultaneously (right axis). In all cases, the 3-cycle coiled PSub successfully causes local stabilization in the flow, with stabilization strength consistent with the corresponding value of the performance metric. The all-modes case demonstrates broadband stabilization spanning the entire unstable region for the given value of ReRe [shown in (a)] with an intensity that matches the ”additive" effect of the $P$ values over the four instability frequencies [shown in (b)]. (d) $K_{\text{p}}$ plot for the 700-Hz instability case showing contrast in performance when the 3-cycle coiled Psub is installed versus using the nominal uncoiled PSub. The results confirm the transformation from destabilization to stabilization due to the coiling.
@@
The results shown in Fig. 2 are obtained by FE analysis, separate from the subsequent fluid-structure simulations we discuss later. The coiled PSub turning junctions at which the rotational locking is applied are denoted Junctions $i=\{1,2,3$,10},10\} in Fig. 1b. The dispersion curves for longitudinal motion for the resonant elastic PSub as a wave propagation medium as well as the frequency response of its finite structural version are calculated and shown in Fig. 2a. We again emphasize that the dispersion curves are immune to the coiling because of the rotational locking [43]. The longitudinal response of the coiled PSub is denoted $\eta^{(i)}$. When analyzing the coiled PSub as a standalone structure, uncoupled to the flow, a complex-valued frequency response function is obtained (see Appendix). This quantity contains information about the displacement amplitude $|\eta^{(i)}|$ and phase $\phi^{(i)}$ of the structure’s response at each junction $i$. In the case of an uncoiled structure (0 coiling cycles), the harmonic excitation is applied at the first junctions ($i=1$) and the complex response is likewise measured at the first junction. For all subsequent coiling configurations (e.g., 1, 2, and 3 coiling cycles), the harmonic excitation is distributed uniformly among the junctions that will be interacting with the flow upon installation, thus its magnitude is divided by the number of junctions excited. For instance, in the case of three coiling cycles, the junctions 1, 3, 5, and 7 are each excited so the magnitude of the harmonic force is $F/4$ for each of these junctions. To obtain the complex frequency response of the coiled cases, the transfer functions for each combination of excited and responsive junctions are calculated and superimposed. For instance, in the case of only 1 coiling cycle, we excite the structure at Junctions 1 and 3, and register the response also at Junctions 1 and 3; so the set of transfer functions here is (1,1), (1,3), (3,1), (3,3), where the first entry of (*,*) represents the junction of input (excitation) and the second entry represents the junction of output (response). In the case of 3 coiling cycles, 16 transfer functions are computed and superimposed, yielding $\eta_{\text{int}}=\Sigma_{i}\eta^{(i)}$. The amplitude and phase of each coiled case are then calculated by taking the absolute value and angle, respectively, of this complex superposition of transfer functions.

### v4

@@ Abstract
We report the discovery of super resonance—a new regime of resonant behavior in which a mode’s out-of-phase response persists far beyond its classical bandwidth. This effect emerges from a coiled phononic structure composed of a locally resonant elastic metamaterial and architected to support multiple internal energy pathways. These pathways converge at a single structural location, enabling extended modal dominance and significantly broadening the frequency range over which a resonant phase is sustained. We demonstrate by direct numerical simulations the implications of this mechanism in the context of flow instability control, where current approaches are inherently constrained by the characteristically narrow spectral bandwidth of conventional resonances. Using a super-resonant phononic subsurface structure interfacing with a channel flow, we show passive simultaneous suppression of four unstable flow perturbations across a frequency range more than five times wider than that is achievable with a standard resonance in an equivalent uncoiled structure. By enabling broadband, passive control of flow instabilities, super resonance overcomes a longstanding limitation in laminar flow control strategies. More broadly, it introduces a powerful new tool for phase-engineered wave–matter interaction. The ability to preserve out-of-phase modal response across wide spectral ranges establishes a fundamental advance in the physics of resonance, with far-reaching implications for suppressing fully developed turbulent flows and beyond.
@@ INTRODUCTION
The harmonic oscillator is one of the most fundamental constructs in physics, serving as a building block across classical mechanics, quantum theory, solid-state physics, and wave dynamics [1, 2, 3, 4, 5]. Its resonance behavior—marked by usually a sharply peaked amplitude response and an associated narrow out-of-phase spectral region—underpins our understanding of modal systems, from phonons and photons to acoustic and electromagnetic cavities. In both isolated and coupled forms, harmonic oscillator modes are intrinsically limited in bandwidth; for coupled systems, the phase response is tightly overridden by the presence of nearby resonances or anti-resonances. This long-standing narrowband constraint has limited the scope of resonance-based control strategies across physics and engineering. In this work, we introduce super resonance: a new regime of modal behavior in which a structural mode sustains an out-of-phase response that persists far beyond its classical resonance bandwidth, despite the nominal presence of adjacent resonant peaks or anti-peaks. This behavior is achieved through a uniquely architected coiled phononic structure featuring multiple internal elastic energy pathways that converge to an effectively single spatial location. This configuration enables extended modal phase persistence, laying the foundation for broadband control in systems traditionally constrained by narrowband resonance dynamics—including the passive suppression of multiple unstable flow perturbations, which is the focus of this investigation.
@@
To overcome these limitations, the concept of phononic subsurfaces (PSubs) has been introduced as a passive intervention approach capable of precise and responsive control of flow instabilities [19]. A PSub comprises a finite phononic structure that is placed beneath the fluid-structure interface. Given its finite size, it represents a truncated [20, 21, 22] phononic material [23, 24, 25, 26], such as a phononic crystal that exhibits Bragg scattering [19, 27, 28, 29, 30], or a locally resonant elastic metamaterial that exhibits resonance hybridizations [28, 31]. The unit-cell and finite-structure characteristics of a PSub may be designed to passively force incident flow instability waves to encounter out-of-phase conditions. This causes a decrease, near the surface, in the production rate of the kinetic energy of the continuously incoming flow perturbation waves, and subsequently their attenuation across the region covered by the PSub. The outcome is a local reduction in skin-friction drag where the PSub unit is installed [31]. Recent developments extended the concept to downstream control, by either adopting a multi-input, multi-output configuration [32], employing a lattice of PSubs [33], or using PSubs tuned to intermediate phase values [34]. The tuning of a single PSub structure depends on knowledge of the frequencies and wavenumbers of the relevant flow instabilities, including incidents involving multiple coexisting modes.
@@
While this passive phonon-enabled approach has shown great promise, in areas ranging from transition delay [32, 33] to control of unsteady flows [35], and recently extending to disturbance attenuation in laminar hypersonic flows [34], its effectiveness remains limited by the characteristically narrowband nature of the modal response in the interfacing structure. This limitation also applies to other conventional resonators such as Helmholtz resonators [36], and other wall-based flow control approaches suffer the same constraint [37]. Extensive investigations have been done on various PSub configurations, but due to this constraint are all confined to spectrally narrow phase response—whether, for example, due to a truncation resonance  [^1] A band-gap truncation resonance may exhibit a modestly broader out-of-phase bandwidth than a standard structural resonance—particularly when the band gap is wide [28]. Yet it remains fundamentally a conventional resonance subject to the same narrowband limitations described above. located within a band gap [19, 27, 29] or a subwavelength structural resonance influenced by the presence of local resonators [28, 31]. This narrowband restriction reduces the applicability of PSubs to realistic settings, such as aircraft flight, where flow instabilities emerge over a relatively broad frequency range  [^2] These instability bands depend on the Reynolds number $Re$ and can be predicted a priori through linear stability analysis [72, 73, 27]. In this paper, by introducing super resonance, we present a new PSub concept that facilitates a significant broadening of the frequency range of operation.
@@
Numerous studies across the physical sciences have investigated different forms of enhanced resonant behavior, often highlighting superior properties. Examples include acoustic super-scattering from elastic shells [40], strong plasmonic field enhancement and lasing [41], high-Q supercavity modes enabled by bound states in the continuum in subwavelength dielectric resonators [42], and intensified Mie resonances in dielectric microspheres [43]. In each case, the response is intensified through distinct means—boundary-induced coherence, material gain, non-radiating interference effects, or structural symmetry, respectively. However, the underlying mechanisms remain rooted in conventional resonant behavior: a narrowband spectral peak that fundamentally constrains modal reach across the frequency domain. In contrast, the current work targets broadening the phase bandwidth of a target resonance beyond its natural limits. The core premise is to intervene in key energy pathways that govern the mechanistic formation of the resonance as it is observed at the response location.
@@
In any resonant system, a peak response arises due to an excitation at a specific spatial location (or domain), and, similarly, there is a distinct spatial location (or domain) where the response may be probed. For a grounded simple harmonic oscillator, there is only a single point of excitation and a single response point, creating a resonant peak with its characteristic narrow bandwidth  [^3] The width of a resonant peak may be broadened by adding dissipation, but at the expense of the response amplitude. Such characteristic narrow band also holds for modes in a coupled oscillator system. Here we consider a multi-degree-of-freedom system composed of coupled oscillators. But we reconfigure the spatial topology to allow multiple pathways for the energy within the structure to converge towards an effectively single response point in space, selected to coincide with the excitation point. In this manner, the modal response of the system exhibits an effectively reconstructed mode with a significantly widened bandwidth, exceeding the fundamental bandwidth associated with a conventional resonant mode.
@@
For application to the flow stabilization problem, the concept is realized in the form of a PSub architecture composed of a coiled beam structure, configured such that more than one location along the structure interfaces with a single point in the flow. The coiling is done by a series of full $180^{\circ}$ turns, incorporating rotational locking at the turning junctions-an approach that fully preserves the phonon band structure compared to the uncoiled configuration [45]. By following this design strategy, we observe remarkable broadening of a performance metric that governs the strength and spectral extent of intervention with the flow [19, 28, 31].
@@
Figure 1: Super-resonant coiled PSub design: A schematic of the concept of a coiled phononic subsurface with multiple structural connectivity. The physical model, shown in (a), features a central voided beam that is rotationally locked at the turning junctions, leading to a full preservation of the phonon band structure [45]. The beam “trunk" has an array of branching resonators in the form of small cantilevered beams-this configuration in its entirety is amenable to fabrication. The dimensions of the PSub unit cell are as follows based on the parameters shown in (a): $L_{\text{UC}}=20$ $m$, $w_{\text{c}}=1.44$ $m$, $h_{\text{c}}=4.63$ $m$, $l_{\text{h}}=1.17$ $m$, $w_{\text{h}}=1.12$ $m$, $l_{1}=18.23$ $m$, $w_{1}=2.06$ $m$, $l_{2}=12.93$ $m$, $w_{2}=4.68$ $m$, $w_{\text{f}}=0.4$ $m$. The central beam component is composed of ABS polymer with density $\rho_{\text{ABS}}=1200$ kg/m3 and Young’s modulus $E_{\text{ABS}}=1$ GPa. The resonant branches are made from aluminium with density $\rho_{\text{Al}}=2700$ kg/m3 and Young’s modulus $E_{\text{Al}}=68.8$ GPa. In our computational investigation, this 3D structure is modeled in the form of a 1D rod, as shown in (b), with effective Young’s modulus $E_{\text{eff}}=0.2$ GPa and density $\rho_{\text{eff}}=450$ kg/m3, containing an array of spring-mass resonators attached based on the following parameters: $L_{\text{UC}}=20$ $m$, $k_{\text{res},1}=2.6\times 10^{10}$ N/m, $k_{\text{res},2}=7.11\times 10^{11}$ N/m, $m_{\text{res},1}=1.62$ g, $m_{\text{res},2}=4.2$ g. Material damping is introduced to the rod component in the form of viscous proportional damping with constants $q_{1}=0$ and $q_{2}=6\times 10^{-8}$ [28]. Given that our interest is in only the longitudinal motion, the rod model provides a good approximation of the coiled PSub response, owing to the rotational locking applied at the junctions. The selected simple model parameters deviate less than 2% from the actual unit-cell equivalent configuration. With one coiling cycle, we have two structural points that are simultaneously interfacing with the flow, labeled Junctions 1 and 3. With two coiling cycles, Junctions 1, 3, and 5 are in contact with the flow, and so on in an odd-number progression. We find that with three coiling cycles, comprising Junctions 1, 3, 5, 7, super-resonant behavior emerges (see Fig. 2). While the coiled PSub design is confined to a specific finite dimension along the spanwise direction, its response can still be considered one-dimensional when normalized by the surface area interacting with the flow, marked in red and denoted by AA. The excitation of the structure is marked by $F$ and the corresponding input into the flow is marked by $\Sigma\eta_{i}$. In this work, we consider only a single coiled PSub covering the full span of the channel. However, application to longer spanwise distances (e.g., for wider channels or boundary-layer flows) may be realized by installation of a periodic arrangement of the coiled PSub along that direction (side view). Additional rows may also be added along the streamwise direction to form a PSub lattice setup [33].
@@
Figure 2: Evidence of super resonance: Coiled PSub dispersion and vibration characteristics: (a) Dispersion, (b-c) harmonic frequency response, and (d) performance metric for the coiled PSub with multiple structural connectivity. The dispersion is unchanged with the number of coiling cycles since the design of the unit cell is the same in all cases and band structure preservation is maintained due to the rotational lockings [45]. The amplitude and phase frequency dependency for the uncoiled (0 coils) and 3-cycle coiled configurations are shown in (b-c). Positive (red) and negative (green) regions of $P$ represent destabilization and stabilization, respectively, upon employment of the PSub for flow control. Zero coiling cycles represent a PSub with 1 point-of-contact with the flow, while 3 coiling cycles represents the same PSub but with 4 points-of-contact with the flow, as demonstraed in Fig. 1. As the number of coils is increased from 1 to 3, the destabilization and stabilization regions across the frequency domain are narrowed and broadened, respectively. With 3 coiling cycles, super resonance emerges yielding a perfectly contiguous stabilization band that is nearly two times larger than that of the uncoiled reference case. A quasi-super resonance regime also emerges, one that is effectively contiguous and nearly four times larger than that of the reference configuration.
@@
In its nominal uncoiled form, the PSub considered comprises a finite linear elastic metamaterial consisting of 10 unit cells, where each is formed from three central voided beam-like segments with two distinct branching cantilever beams in-between, serving as local resonators; see illustration with geometric details in Fig. 1. To make extensive coupled fluid-structure simulations feasible, this phononic structure is modeled as an elastic rod with two mass-spring resonators attached per unit cell, similar to the configuration of Ref. [28]. The underlying reduce-order-modeling is facilitated by a combined homogenization and machine learning-algorithm [55, 56]; see Appendix for details. The PSub is free to deform at the edge interfacing with the flow (top) as well as all other intermediate locations throughout its structure, but is fixed at the other end (bottom). The PSub elastic domain is allowed to deform in complete independence from the adjacent rigid wall, thus the flow “experiences" a deformation across a given fluid-PSub interface junction [19, 28]. The length of the unit cell along the wall-normal direction is $L_{\mathrm{UC}}=2$ $\mathrm{cm}$ (i.e., total uncoiled PSub length is 2020 $\mathrm{cm}$). The resonator frequencies are set to $\Omega_{\mathrm{res,1}}=5000$ $\mathrm{Hz}$ and $\Omega_{\mathrm{res,2}}=20,000$ $\mathrm{Hz}$ by tuning their point masses to be three and five times heavier than the total mass of the unit-cell base, i.e., $m_{\mathrm{res,1}}=3\times\rho L_{\mathrm{UC}}$ and $m_{\mathrm{res,2}}=5\times\rho L_{\mathrm{UC}}$, where $\rho$ is the base material effective density. Hence, the stiffnesses of the resonator springs are $k_{\mathrm{res,n}}=m_{\mathrm{res,n}}(2\pi f_{\mathrm{res,n}})^{2}$, where $n=1,2$.
@@
 The PSub coiling protocol is done as follows. A full $180^{\circ}$ turn is applied to each unit cell, starting with the second one, and up to twice the number of coiling cycles. Thus, with $n$ coiling cycles, $2n$ full turns are applied at the intersections of the first $2n+1$ unit cells. In order to guarantee that the phonon band structure is preserved, rotationally rigid junctions are placed at these intersections to lock the rotation degree of freedom [45]. On the top side, these junctions converge to a $n+1$ points in contact with the flow, with the ability to interact independently with it. At these points, the flow represents the excitation source from the view point of the coiled PSub. Since these points are confined within a small control region along the streamwise direction compared to the perturbation wavelength(s), from the flow point of view they are effectively perceived as a single response point. This means that the rotationally rigid junctions transmit and distribute the total force equally from the flow to the coiled PSub, and the coiled PSub’s response is the cumulative contribution of all the points in contact with the flow.In this manner, the fluid-structure interaction takes a multiple input, multiple output form for the PSub and a single input, single output form for the flow. See Fig. 1 for a schematic and geometric details.
@@
Figure 3: Super-resonant coiled PSub performance by DNS: (a) Stability map in the frequency-Reynolds number domain obtained by solving the Orr-Sommerfeld equation [47, 48] for the channel. (b) Zoom-in of the performance metric plot for the super-resonant 3-cycle coiled PSub, with frequencies of the four unstable TS waves marked. (c) Time-averaged perturbation kinetic energy in the flow as a function of streamwise position in the channel. The results of five separate simulations are shown: four in which a distinct TS instability wave is controlled (left axis), and one where all four TS instability waves are controlled simultaneously (right axis). In all cases, the 3-cycle coiled PSub successfully causes local stabilization in the flow, with stabilization strength consistent with the corresponding value of the performance metric. The all-modes case demonstrates broadband stabilization spanning the entire unstable region for the given value of ReRe [shown in (a)] with an intensity that matches the ”additive" effect of the $P$ values over the four instability frequencies [shown in (b)]. (d) $K_{\text{p}}$ plot for the 700-Hz instability case showing contrast in performance when the 3-cycle coiled Psub is installed versus using the nominal uncoiled PSub. The results confirm the transformation from destabilization to stabilization due to the coiling.
@@
The results shown in Fig. 2 are obtained by FE analysis, separate from the subsequent fluid-structure simulations we discuss later. The coiled PSub turning junctions at which the rotational locking is applied are denoted Junctions $i=\{1,2,3$,10},10\} in Fig. 1b. The dispersion curves for longitudinal motion for the resonant elastic PSub as a wave propagation medium as well as the frequency response of its finite structural version are calculated and shown in Fig. 2a. We again emphasize that the dispersion curves are immune to the coiling because of the rotational locking [45]. The longitudinal response at a given junction in the coiled PSub is denoted $\eta_{i}$. When analyzing the coiled PSub as a standalone structure, uncoupled to the flow, a complex-valued frequency response function is obtained (see Appendix). This quantity contains information about the displacement amplitude $|\eta_{i}|$ and phase $\phi_{i}$ of the structure’s response at each junction $i$. In the case of an uncoiled structure (0 coiling cycles), the harmonic excitation is applied at the first junctions ($i=1$) and the complex response is likewise measured at the first junction. For all subsequent coiling configurations (e.g., 1, 2, and 3 coiling cycles), the harmonic excitation is distributed uniformly among the junctions that will be interacting with the flow upon installation, thus its magnitude is divided by the number of junctions excited. For instance, in the case of three coiling cycles, the junctions 1, 3, 5, and 7 are each excited so the magnitude of the harmonic force is $F/4$ for each of these junctions. To obtain the complex frequency response of the coiled cases, the transfer functions for each combination of excited and responsive junctions are calculated and superimposed. For instance, in the case of only 1 coiling cycle, we excite the structure at Junctions 1 and 3, and register the response also at Junctions 1 and 3; so the set of transfer functions here is (1,1), (1,3), (3,1), (3,3), where the first entry of (*,*) represents the junction of input (excitation) and the second entry represents the junction of output (response). In the case of 3 coiling cycles, 16 transfer functions are computed and superimposed, yielding $\eta_{\text{int}}=\Sigma_{i}\eta_{i}$. The amplitude and phase of each coiled case are then calculated by taking the absolute value and angle, respectively, of this complex superposition of transfer functions.